# Mitigation of multi-path propagation artefacts in acoustic targets with cepstral adaptive filtering

### v1

@@ IV-B Underwater Acoustic Signal Classification
The sound emitted by a ship is composed of machinery and propeller noises, which appear as constant lower frequencies components in the frequency spectra, and hydrodynamic noise, which is a broadband component that shifts between mid and high frequencies according to the ship’s speed [carlton2018marine]. Considering ships as moving targets, the propagation of their radiated noise in the ocean is also affected by the LME. However, in this case, the effect contains valuable information about the depth of the source, which is intrinsically related to the ships’ physical structure. To account for that, the LME is used as a complementary information in the classification pipeline, which considers the original audio together with the LME filtered version as inputs to the classification, as illustrated in Figure 5.
@@
A common strategy in the ship-type classification literature is to apply deep learning algorithms in spectrogram-like representations [domingos2022survey]. Knowing that the ship radiated noise is most present on the lower frequency spectra, the CQT is used to obtain the spectrogram. The time-filtering of the cepstral components was also performed in the CQT, aiming to evaluate the transferability of the strategy to other types of spectrogram and applications. As the CQT presents a logarithmic scale rather than a linear one, the cepstral representation obtained is not in the time domain as quefrencies, thus the components are denoted as cepstral coefficients.
@@ IV-B1 Implementation Details
The machine learning pipeline used for classification was adapted from [domingos2025improving], which employs the GSE ResNeXt model due to its stability and fast convergence. To maintain a consistent training environment, the model was trained using the AdamW optimiser with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$ for regularisation. Training was conducted over 50 epochs with a mini-batch size of 64.
For deep learning classification tasks, dataset selection plays a crucial role, as models tend to replicate patterns and trends present in the training data. Additionally, large-scale datasets are typically required to ensure model convergence. Considering this need, the Ocean Networks Canada Society database [^1] https://oceannetworks.ca, a publicly available source of unlabelled underwater audio recordings, was selected. Due to the availability of Automatic Identification System (AIS) data for annotation, three time windows recorded between May 2016 and October 2018 at the Strait of Georgia are commonly used in the literature [irfan2021deepship, domingos2022investigation]. These recordings were acquired using an icListen smart hydrophone, with a bandwidth of 1Hz-12kHz and a sampling rate of 32, [^0] Hz. The sensors were deployed at depths of 141m, 147m and 144m for each recording window, representing a shallow-water scenario with favourable conditions for LME occurrence. To expand the scope of the analysis and enable performance comparisons, the following two subsets of this data were used as reference:
@@
Contains recordings annotated into four classes based on vessel type (tug, passenger ship, cargo ship, and oil tanker) [irfan2021deepship]. To avoid interference from overlapping ship signals, only recordings from a single vessel within a 2 km radius of the hydrophone were included in the database. If a second vessel entered this radius, the recording was truncated and the remainder discarded. The dataset comprises 613 recordings from 265 distinct ships, totalling approximately 47 hours of audio.
@@
Using an automatic annotation pipeline [domingos2022investigation], recordings from ten ship-type classes were collected: tug boat, passenger ship, cargo ship, oil tanker, pilot vessel, rescue, sailing, pleasure craft, fishing, and dredger. Unlike DeepShip, this dataset includes recordings from vessels within a 2 km radius of the hydrophone, under the condition that no other vessel is present within a 4 km radius. This constraint was applied to minimise interference from nearby vessels. The dataset contains 1,111 recordings from 479 different ships, with a total duration of approximately 116 hours.
@@ IV-B2 Dataset Segmentation Tasks
In underwater acoustic classification, dataset preprocessing typically involves segmenting full-length audio recordings into smaller parts, which can then be used independently during training, testing, or validation stages. Three segmentation strategies are commonly applied [domingos2025improving, liu2021underwater]: Task 1 follows a segment-first approach, where the full audio is first segmented, the segments are randomised, and then divided into training and testing subsets; Task 2, referred to as Sequential Audio Sampling (SAS), assigns consecutive portions of each audio file to either training or testing, followed by segmentation into smaller parts; Task 3, referred to as Full-Length Audio Sampling (FAS), uses each full audio file as either a training or testing sample, which is then segmented into smaller parts. Due to the similarity between adjacent recordings, Task 1 is typically prone to overfitting [domingos2025improving]. Therefore, in this study, only Task 2 and Task 3 were considered, and are referred to as SAS and FAS, respectively.
@@ IV-B3 Classification Metrics
In this research, classification performance is evaluated using the Matthews Correlation Coefficient (MCC) [matthews1975comparison]. The MCC is particularly suitable for handling class imbalances in classification tasks, as it assigns a high score to predictions that are accurate across all four base categories (true positives ($tp$), true negatives ($tn$), false positives ($fp$) and false negatives ($fn$)). The coefficient is defined as:
@@
Given that significant variations in classification performance are commonly observed across folds in underwater acoustic classification [domingos2025improving], this research reports the MCC alongside the difference from the baseline ($\Delta$). This additional information enables direct comparison of the results with the baselines, while also accounting the impact of standard deviation in assessing statistical relevance, thereby reducing inconsistencies across folds.
@@ V-A Multi-path Reflection Attenuation in Synthetic Data
Considering the dataset of acoustic data with simulated movement, the cepstral time-filtering was performed in the spectrograms using the adaptive band-stop filter. The filtering was applied in both the acoustic spectrographic data considering only the direct path component, here denoted as $A(f,t)$, and the data considering the waves propagated trough direct and reflected paths, denoted as $B(f,t)$. Using the spectrogram of the acoustic data with amplitude attenuation as reference ($R(f,t)$), the similarity between the original signal and the reference was compared against the similarity between the filtered signal ($A^{\prime}(f,t)$ and $B^{\prime}(f,t)$) and the reference. Figure 7 shows an example data containing the spectrograms and cepstrograms of the reference, delayed signals, and filtered acoustic signals.
Figure 7: Comparison of the spectrograms (upper row) and cepstrograms (bottom row) of the synthetic dataset components filtered using the time filtering of the cepstrum.
@@
The spectrograms containing the delay caused by the direct path of wave propagation ($A(f,t)$) exhibited an S-shape curve localised at the simulated closest point of approach (see the frequency decaying around 5 seconds in Figure 7), posing as a challenge for the filtering algorithm due to the frequency shift caused by Doppler Effect. Figure 8 shows the behaviour of the similarity and distortion metrics across all velocities for both $R(f,t),A(f,t)$ and $R(f,t),A^{\prime}(f,t)$ pairs. As it is more sensitive to the absolute values, the SNR measurement showed a decrease in the ratio of signal of the filtered version in comparison with the unfiltered, as can be seen in graph (a). However, when considering the logarithmically compressed spectrogram in the LSD and IS distortion metrics, the filtering procedure demonstrated beneficial effects, as indicated by the lower distortion values obtained for the $R(f,t),A^{\prime}(f,t)$ pair in the graphs shown in Figure 8 (b) and (c). As they reflect the perceptual quality of the audio, the filtering process was successful in improving the quality of signals affected by the Doppler effect. As expected, the quality of both pairs decrease as the velocity increases, which reflects the intensity of the synthetic movement effects. The IS distance metric points that the filtering procedure is exceptionally useful with increasing velocities, showing a distancing trend of both curves in Figure 8 (c), with distortion reductions ranging from 15% to 88%.
Figure 8: Similarity metrics obtained during the comparison of the reference $R(f,t)$ with the signal propagated over direct path ($R(f,t)\times A(f,t)$) and the filtered version ($R(f,t)\times A^{\prime}(f,t)$) over different velocities. SNR is represented in (a), LSD in (b), and IS in (c).
@@
Discrepancies between the SNR measurements and distortion metrics, such as LSD and IS, are related to the amplitude of the filtered signals. As SNR is sensitive to absolute values of magnitude, variations in the amplitude obtained after filtering have a substantial impact in the final metric [gray1980distortion]. As $A(f,t)$ only contains the directly propagated signal, as well as $R(f,t)$, the SNR between both will be higher when compared with the filtered version, which can suffer from amplitude attenuation. For the distortion metrics, which consider the ratio of the log components, the measurements are lower after the filtering process as the spectral shape is recovered.
@@
The evaluation of the spectrogram containing both the waves propagated in the direct and reflected paths $B(f,t)$ and its filtered version $B^{\prime}(f,t)$ exhibited positive results for the filtering strategy. As can be seen in Figure 9, the comparison of the metrics curves between the two pairs of representations, $R(f,t),B(f,t)$ and $R(f,t),B^{\prime}(f,t)$, showed a consistent quality improvement after the filtering process for all velocities. As the artefacts presented in $B(f,t)$ contemplate not only the Doppler effect but also the LME, the impact of the filtering strategy was more prominent, reflecting in all the three metrics, as can be seen in Figure 9 curves. Despite the increased distance between the curves in graphs (b) and (c), the behaviour of the LSD and IS metrics was similar to that in the previous experiment, reflecting the improvements obtained after filtering. Also, graph (a) shows that the SNR of the filtered version was higher than the unfiltered, reflecting the capability of the filtering process of removing the echoes caused by the reflection components.
Figure 9: Similarity metrics obtained during the comparison of the reference $R(f,t)$ with the signal propagated over direct and reflected paths ($R(f,t)\times B(f,t)$) and the filtered version ($R(f,t)\times B^{\prime}(f,t)$) over different velocities. SNR is represented in (a), LSD in (b), and IS in (c).
The SNR values obtained for the second experiment (as can be seen in graph (a) from Figure 9) are higher after filtering due to the recovering of the original sound, which removes the reflections. As the reflections cause a constructive and destructive pattern which will also change the absolute amplitude values, their removal will have a great impact in the SNR when comparing with a reference that does not contain the reflection components. For both the LSD and IS metrics, once again, the spectral shape was recovered after the filtering, showing a perceptual improvement of the signal. The IS measure in graph (c), specifically, point out to an increasing importance of the filtering for higher velocities, as the curves show a distancing trend for both experiments, as can be seen in Figure 9 and 8. The increase in LSD and IS metrics also point to a perceptual quality improvement after filtering application, which can be beneficial for recognition and classification algorithms which take the spectrograms as input.
@@ V-B LME Filtering in Underwater Acoustic Classification
The filtering method was applied to the CQT cepstral coefficients of the underwater acoustic data to attenuate the LME present in CQT spectrograms. Figure 10 shows the qualitative attenuation of LME and movement artefacts after the filtering process. In the CQT spectrogram of the unprocessed signal, the LME effect is visible at lower frequencies, resulting from reflections of the sound wave on the seafloor. This pattern is also evident in the cepstral coefficients, appearing as diagonal lines in the intermediate coefficients. The behaviour of the CQT cepstral coefficients is analogous to that observed in cepstrograms, confirming that reflection patterns are identifiable in this representation.
Figure 10: Comparison of the CQT spectrograms (top row) and cepstrograms (bottom row) of the unprocessed and filtered representations. In the unprocessed signal, the u-shaped patterns are caused by the LME. After processing, only the straight horizontal lines are maintained, which represent the harmonic components of the ship-radiated noise.
After filtering, both the unprocessed and filtered signals were used as inputs to a classification pipeline. The results, presented in Table I, show the performance of the classification algorithm across three scenarios: A classification pipeline using the unprocessed spectrogram alone; a pipeline which uses the spectrogram after the time-filtering of the cepstrum; and the combination of the unprocessed and the filtered spectrograms in a parallel classification, as illustrated in Figure 5.
@@
Filtered alone 77.65 ± 2.46 - 0.44 ± 3.14 65.56 ± 5.08 - 0.89 ± 0.95
Unprocessed + Filtered 80.37 ± 2.70 + 2.28 ± 1.30 67.79 ± 5.19 + 1.35 ± 0.54
@@
Filtered alone 79.26 ± 1.49 - 0.45 ± 0.69 50.59 ± 0.62 - 0.92 ± 0.96
Unprocessed + Filtered 82.33 ± 1.71 + 2.62 ± 0.54 52.36 ± 0.30 + 0.85 ± 0.50
The results indicate that using the filtered representation alone does not improve classification performance for SAS and FAS on both datasets. Given that the average differences ($\Delta$) are smaller than their respective standard deviations, the filtered representation alone does not yield consistent improvements. However, the parallel filtering pipeline, which combines both filtered and unprocessed representations, demonstrates a noticeable improvement in classification performance across both tasks and datasets. SAS showed the most significant gains, with increases of 3.46% and 3.29% for DeepShip and VTUAD v2, respectively. FAS also benefited from the parallel approach, with improvements of 2.03% and 1.65% for DeepShip and VTUAD v2, respectively.
As expected, the use of filtering alone was not beneficial. The filtering strategy proved effective in removing reflected components from the audio, removing also the information encoded in these components, such as the sound source depth. This information may be intrinsically linked to ship size and tonnage, which are important features for distinguishing between vessel types. Therefore, although the filter successfully mitigates the influence of multi-path reflections, it may also remove discriminative features that are valuable for classification.
Since SAS utilises different segments of the same audio for training and testing, the attenuation of time-varying components proved beneficial for classification. In contrast, FAS benefited more from the complementary information provided by the LME. Since ship-radiated noise spectrogram contains hydrodynamic components in the form of broadband signals, the LME present in these signals encodes information about the physical structure of the ship. Retaining this information in the unprocessed signal contributed positively to classification. However, as the engine and machinery sounds were hidden in the reflection artefacts, the filtering strategy contributed to the identification of harmonic patterns present in the spectrograms. In practice, the parallel classification pipeline provides both representations separately to the model, offering complementary perspectives that enhance performance.

### v2

@@ IV-B Underwater Acoustic Signal Classification
The sound emitted by a ship is composed of machinery and propeller noises, which appear as constant lower frequencies components in the frequency spectra, and hydrodynamic noise, which is a broadband component that shifts between mid and high frequencies according to the ship’s speed [8]. Considering ships as moving targets, the propagation of their radiated noise in the ocean is also affected by the LME. However, in this case, the effect contains valuable information about the depth of the source, which is intrinsically related to the ships’ physical structure. To account for that, the LME is used as a complementary information in the classification pipeline, which considers the original audio together with the LME filtered version as inputs to the classification, as illustrated in Figure 5.
@@
A common strategy in the ship-type classification literature is to apply deep learning algorithms in spectrogram-like representations [10]. Knowing that the ship radiated noise is most present on the lower frequency spectra, the CQT is used to obtain the spectrogram. The time-filtering of the cepstral components was also performed in the CQT, aiming to evaluate the transferability of the strategy to other types of spectrogram and applications. As the CQT presents a logarithmic scale rather than a linear one, the cepstral representation obtained is not in the time domain as quefrencies, thus the components are denoted as cepstral coefficients.
@@ IV-B1 Implementation Details
The machine learning pipeline used for classification was adapted from [12], which employs the GSE ResNeXt model due to its stability and fast convergence. To maintain a consistent training environment, the model was trained using the AdamW optimiser with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$ for regularisation. Training was conducted over 50 epochs with a mini-batch size of 64.
For deep learning classification tasks, dataset selection plays a crucial role, as models tend to replicate patterns and trends present in the training data. Additionally, large-scale datasets are typically required to ensure model convergence. Considering this need, the Ocean Networks Canada Society database [^1] https://oceannetworks.ca, a publicly available source of unlabelled underwater audio recordings, was selected. Due to the availability of Automatic Identification System (AIS) data for annotation, three time windows recorded between May 2016 and October 2018 at the Strait of Georgia are commonly used in the literature [22, 11]. These recordings were acquired using an icListen smart hydrophone, with a bandwidth of 1Hz-12kHz and a sampling rate of 32, [^0] Hz. The sensors were deployed at depths of 141m, 147m and 144m for each recording window, representing a shallow-water scenario with favourable conditions for LME occurrence. To expand the scope of the analysis and enable performance comparisons, the following two subsets of this data were used as reference:
@@
Contains recordings annotated into four classes based on vessel type (tug, passenger ship, cargo ship, and oil tanker) [22]. To avoid interference from overlapping ship signals, only recordings from a single vessel within a 2 km radius of the hydrophone were included in the database. If a second vessel entered this radius, the recording was truncated and the remainder discarded. The dataset comprises 613 recordings from 265 distinct ships, totalling approximately 47 hours of audio.
@@
Using an automatic annotation pipeline [11], recordings from ten ship-type classes were collected: tug boat, passenger ship, cargo ship, oil tanker, pilot vessel, rescue, sailing, pleasure craft, fishing, and dredger. Unlike DeepShip, this dataset includes recordings from vessels within a 2 km radius of the hydrophone, under the condition that no other vessel is present within a 4 km radius. This constraint was applied to minimise interference from nearby vessels. The dataset contains 1,111 recordings from 479 different ships, with a total duration of approximately 116 hours.
@@ IV-B2 Dataset Segmentation Tasks
In underwater acoustic classification, dataset preprocessing typically involves segmenting full-length audio recordings into smaller parts, which can then be used independently during training, testing, or validation stages. Three segmentation strategies are commonly applied [12, 28]: Task 1 follows a segment-first approach, where the full audio is first segmented, the segments are randomised, and then divided into training and testing subsets; Task 2, referred to as Sequential Audio Sampling (SAS), assigns consecutive portions of each audio file to either training or testing, followed by segmentation into smaller parts; Task 3, referred to as Full-Length Audio Sampling (FAS), uses each full audio file as either a training or testing sample, which is then segmented into smaller parts. Due to the similarity between adjacent recordings, Task 1 is typically prone to overfitting [12]. Therefore, in this study, only Task 2 and Task 3 were considered, and are referred to as SAS and FAS, respectively.
@@ IV-B3 Classification Metrics
In this research, classification performance is evaluated using the Matthews Correlation Coefficient (MCC) [30]. The MCC is particularly suitable for handling class imbalances in classification tasks, as it assigns a high score to predictions that are accurate across all four base categories (true positives ($tp$), true negatives ($tn$), false positives ($fp$) and false negatives ($fn$)). The coefficient is defined as:
@@
Given that significant variations in classification performance are commonly observed across folds in underwater acoustic classification [12], this research reports the MCC alongside the difference from the baseline ($\Delta$). This additional information enables direct comparison of the results with the baselines, while also accounting the impact of standard deviation in assessing statistical relevance, thereby reducing inconsistencies across folds.
@@ V-A Multi-path Reflection Attenuation in Synthetic Data
Considering the dataset of acoustic data with simulated movement, the cepstral time-filtering was performed in the spectrograms using the adaptive band-stop filter. The filtering was applied in both the acoustic spectrographic data considering only the direct path component, here denoted as $A(f,t)$, and the data considering the waves propagated trough direct and reflected paths, denoted as $B(f,t)$. Using the spectrogram of the acoustic data with amplitude attenuation as reference ($R(f,t)$), the similarity between the original signal and the reference was compared against the similarity between the filtered signals ($A^{\prime}(f,t)$ and $B^{\prime}(f,t)$) and the reference. Figure 7 shows an example data containing the spectrograms and cepstrograms of the reference, delayed signals, and filtered acoustic signals.
Figure 7: Comparison of the spectrograms (upper row) and cepstrograms (bottom row) of the synthetic dataset components, filtered using temporal cepstral filtering. Reflection components are visible in the cepstrograms of $A(f,t)$ (at quefrencies below 1 ms) and $B(f,t)$ (at quefrencies between 1 and 2 ms), in contrast to $R(f,t)$. After filtering, these artefacts are attenuated in the cepstrograms of both $A^{\prime}(f,t)$ and $B^{\prime}(f,t)$, with the attenuation also observable in the corresponding recovered spectrograms.
@@
The spectrograms containing the delay caused by the direct path of wave propagation ($A(f,t)$) exhibited an S-shape curve localised at the simulated closest point of approach (see the frequency decaying around 5 seconds in Figure 7), posing as a challenge for the filtering algorithm due to the frequency shift caused by Doppler Effect. Figure 8 shows the behaviour of the similarity and distortion metrics across all velocities for both $R(f,t),A(f,t)$ and $R(f,t),A^{\prime}(f,t)$ pairs. As it is more sensitive to the absolute values, the SNR measurement showed no improvement in the ratio of signal of the filtered version in comparison with the unfiltered, as can be seen in graph (a). However, when considering the logarithmically compressed spectrogram in the LSD distortion metric, the filtering procedure demonstrated beneficial effects, as indicated by the lower distortion values obtained for the $R(f,t),A^{\prime}(f,t)$ pair in the graphs shown in Figure 8 (b). As it reflects the perceptual quality of the audio, the filtering process was successful in improving the quality of signals affected by the Doppler effect. As expected, the quality of both pairs decrease as the velocity increases, which reflects the intensity of the synthetic movement effects.
Figure 8: Similarity metrics obtained when comparing the reference signal $R(f,t)$ with a signal propagated along the direct path only ($R(f,t)\times A(f,t)$) and its filtered version ($R(f,t)\times A^{\prime}(f,t)$), across different target velocities. Results are shown as the median value with shaded areas indicating the lower and upper quartiles. While cepstral filtering reduced spectral distortion, as measured by LSD, a reduction in SNR was noted in the absence of multi-path effects.
@@
Discrepancies between the SNR measurements and LSD are related to the amplitude of the filtered signals. As SNR is sensitive to absolute values of magnitude, variations in the amplitude obtained after filtering have a substantial impact in the final metric [21]. As $A(f,t)$ only contains the directly propagated signal, as well as $R(f,t)$, the SNR between both will be similar, or sometimes higher, when compared with the filtered version, which can suffer from amplitude attenuation. For the distortion metric, which consider the ratio of the log components, the measurements are lower after the filtering process as the spectral shape is recovered.
@@
The evaluation of the spectrogram containing both the waves propagated in the direct and reflected paths $B(f,t)$ and its filtered version $B^{\prime}(f,t)$ exhibited positive results for the filtering strategy. As can be seen in Figure 9, the comparison of the metrics curves between the two pairs of representations, $R(f,t),B(f,t)$ and $R(f,t),B^{\prime}(f,t)$, showed a consistent quality improvement after the filtering process across all velocities. As the artefacts presented in $B(f,t)$ contemplate not only the Doppler effect but also the LME, the impact of the filtering strategy was more prominent, reflecting in both metrics, as can be seen in Figure 9 curves. Despite the increased distance between the curves in graph (b), the behaviour of the LSD was similar to that in the previous experiment, reflecting the improvements obtained after filtering. Also, graph (a) shows that the SNR of the filtered version was higher than the unfiltered, reflecting the capability of the filtering process of attenuating the echoes caused by the reflection components.
Figure 9: Similarity metrics obtained when comparing the reference signal $R(f,t)$ with a signal propagated along both direct and reflected paths ($R(f,t)\times B(f,t)$) and its filtered version ($R(f,t)\times B^{\prime}(f,t)$), across different target velocities. Results are shown as the median value with shaded areas indicating the lower and upper quartiles. Cepstral filtering improved both LSD and SNR in the multi-path scenario.
The SNR values obtained in the second experiment had a greater increase after filtering when compared with the first experiment (see graphs (a) on Figure 8 and Figure 9). As reflections introduced constructive and destructive interference that altered the signal amplitude, their removal had a pronounced effect on SNR when comparison is made to a reference that did not contain reflected components. For LSD, the filtering process similarly leads to recovery of the spectral shape, reflecting an improvement in perceptual signal quality.
In addition to the previous experiments, a fixed band‑stop filter with an analogous frequency band was used as comparison to the adaptive band‑stop filter. To facilitate visualisation and comparison, only the absolute difference between the noisy and filtered metrics were considered. The results in Figure 10 are summarised by the median across velocities, with shaded bands indicating the lower and upper quartiles. Both the direct‑path experiment (column (a)) and the multi‑path experiment including LME interactions (column (b)) are reported. In column (a), the median SNR shows modest gains for velocities above 50 m/s with the adaptive filter; however, these gains are not substantial once the quartiles are considered. A similar pattern is observed in column (b), with improvements across all velocities and reduced inter-quartile overlap, indicating a more pronounced improvement in SNR. For LSD, both experiments show improvements with adaptive filtering relative to fixed filtering. The effect is most pronounced in the multi‑path scenario (column (b)), with gains ranging from 4–8% on the median value across velocities with no inter-quartile overlap.
Figure 10: Comparison of metric differences ($\Delta$) between the reference and noisy signals ($R(f,t)\times A(f,t)$ and $R(f,t)\times B(f,t)$) and the corresponding filtered signals ($R(f,t)\times A^{\prime}(f,t)$ and $R(f,t)\times B^{\prime}(f,t)$), for both adaptive and fixed band‑stop filters. Column (a) shows results for the experiment involving propagation along the direct path only ($A(f,t)$), whereas column (b) corresponds to propagation along both direct and reflected paths ($B(f,t)$). Results are presented using the median value, with shaded regions indicating the lower and upper quartiles. Adaptive filtering leads to greater improvements in LSD in both experiments, particularly in the attenuation of multi‑path artefacts.
Due to the dynamic nature of the target signature and the environment, adaptive filtering exhibited a better performance than fixed filtering. In multi‑path reflection scenarios, where signal variation is more pronounced, the benefit of adaptive filtering is most evident. The observed improvements in distortion metrics (e.g., LSD) indicate enhanced perceptual signal quality following filtering. The impact of these improvements on recognition and classification algorithms that operate directly on spectrogram‑based representations will be further evaluated.
@@ V-B LME Filtering in Underwater Acoustic Classification
The filtering method was applied to the CQT cepstral coefficients of the underwater acoustic data to attenuate the LME present in CQT spectrograms. Figure 11 shows the qualitative attenuation of LME and movement artefacts after the filtering process. In the CQT spectrogram of the unprocessed signal, the LME effect is visible at lower frequencies, resulting from reflections of the sound wave on the seafloor. This pattern is also evident in the cepstral coefficients, appearing as diagonal lines in the lower coefficients (around the bin 25 in Figure 11). The behaviour of the CQT cepstral coefficients is analogous to that observed in cepstrograms, confirming that reflection patterns are identifiable in this representation.
Figure 11: Comparison of the CQT spectrograms (top row) and cepstrograms (bottom row) of the unprocessed and filtered representations. In the unprocessed spectrogram signal, the u-shaped patterns are caused by the LME. After processing, only the straight horizontal lines are maintained, which represent the harmonic components of the ship-radiated noise.
After filtering, both the unprocessed and filtered signals were used as inputs to a classification pipeline. The results, presented in Table I, show the performance of the classification algorithm across five scenarios: A classification pipeline using the unprocessed spectrogram alone; a pipeline which uses the spectrogram after the time-filtering of the cepstrum for both adaptive and fixed band-stop filtering; and the combination of the unprocessed and the filtered spectrograms in a parallel classification for both filtering strategies, as illustrated in Figure 5.
@@
Fixed Filtered alone 77.18 ± 2.03 -0.91 ± 2.17 66.08 ± 4.94 -0.37 ± 3.03
Adaptive Filtered alone 77.65 ± 2.46 - 0.44 ± 3.14 65.56 ± 5.08 - 0.89 ± 0.95
Unprocessed + Fixed 79.62 ± 2.86 + 1.53 ± 1.25 68.16 ± 4.74 + 1.71 ± 1.04
Unprocessed + Adaptive 80.37 ± 2.70 + 2.28 ± 1.30 67.79 ± 5.19 + 1.35 ± 0.54
@@
Fixed Filtering 79.80 ± 1.28 0.09 ± 1.09 50.03 ± 0.99 -1.47 ± 0.96
Adaptive Filtered alone 79.26 ± 1.49 - 0.45 ± 0.69 50.59 ± 0.62 - 0.92 ± 0.96
Unprocessed + Fixed 82.55 ± 1.32 + 2.84 ± 0.74 52.23 ± 0.83 + 0.72 ± 0.34
Unprocessed + Adaptive 82.33 ± 1.71 + 2.62 ± 0.54 52.36 ± 0.30 + 0.85 ± 0.50
The results indicate that using the filtered representation alone does not improve classification performance for SAS and FAS on either dataset. Consistent with this, the average differences ($\Delta$) are smaller than their respective standard deviations, indicating that the filtered representation alone does not yield consistent improvements. On the other hand, the parallel filtering pipeline, which combines filtered and unprocessed representations, yields noticeable improvements in classification performance across both tasks and datasets. SAS showed the most significant gains, with increases of 3.46% and 3.29% for DeepShip and VTUAD v2, respectively, when using the adaptive filtering. FAS also benefited from the parallel approach, with improvements of 2.03% and 1.65% for DeepShip and VTUAD v2, respectively. The differences between fixed and adaptive filtering are generally subtle, often within one standard deviation. An exception occurs for SAS under the parallel pipeline, where adaptive filtering surpasses fixed filtering by a small margin.
As expected, applying filtering in isolation was not beneficial. The filtering strategy proved effective in attenuating reflected components; however, it also suppressed information encoded in those components, such as the sound source depth and range [16]. This information can correlate with target characteristics (e.g., ship length, tonnage) and thus with vessel type, making them potentially discriminative [40]. Therefore, while the filter successfully mitigates the influence of multi-path reflections, it may also remove discriminative features that are valuable for classification.
Since SAS utilises different segments of the same audio for training and testing, the attenuation of time-varying components proved beneficial for classification. In contrast, FAS benefited more from the complementary information provided by the LME. Because the acoustic source in a ship is related to its physical dimensions, such as the draught and the propeller position, the LME pattern indirectly provides information about the ship’s size and draft[40, 24]. Retaining this information in the unprocessed signal contributed positively to classification. However, as the engine and machinery sounds were hidden in the reflection artefacts [3], the filtering strategy contributed to the identification of harmonic patterns present in the spectrograms. In practice, the parallel classification pipeline provides both representations separately to the model, offering complementary perspectives that enhance performance.